Below are data from a regular survey of about 5000 households, called the Current Population Survey, which are grouped into different demographic categories.

Demographic Group

No. of people

Housewives (age: above 20)

500,700

Full time students (age: above 16)

960,500

Full time workers (age: 20-65)

2,000,600

Part time workers (age: 10-15)

500,000

Part time workers (age: 19-35)

1,750,000

Employed in father’s business (age: above 18)

300,000

Workers (age: 17-60)- temporary on leave

75,000

Workers(age: 17-60)- waiting to be recalled to their old job

40,000

Does not have job but actively looking for job for past 6 weeks  (age: above 16)

90,000

Elderly people who cannot work (age: above 65)

64,000

Use the information to calculate:


a) Adult population.

b) Labor force.

c) Not in the labor force.

d) Labor force participation rate.

e) Unemployment rate.



1
Expert's answer
2021-01-12T13:37:22-0500

a) Adult population.

2000600+1750000+90000+64000+40000+960500=4,905,100

b) Labor force.

2000600+1750000+40000+64000=3,944,600

c) Not in the labor force.

64000+40000=104,000

d) Labor force participation rate.

3944600/4905100=0.80

e) Unemployment rate.

154000/3944600*100=3.90